# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

### v1

@@ Abstract
This paper develops a unified framework of high-order accumulative regularization (AR) framework for convex and uniformly convex gradient norm minimization. Existing high-order methods often exhibit a gap: the function value residual decreases fast, while the gradient norm converges much slower. To close this gap, we introduce AR that systematically transforms fast function value residual convergence rate into fast (matching) gradient norm convergence rate.
Specifically, for composite convex problems, for computing an approximate solution such that the norm of its (sub)gradient does not exceed $\varepsilon,$ the proposed AR methods match the best corresponding convergence rate for the function value residual. We further extend the framework to uniformly convex settings, establishing linear, superlinear and sublinear convergence of the gradient norm under different lower curvature conditions. Moreover, we design parameter-free algorithms that require no input of problem parameters, e.g., Lipschitz constant of the $p$-th order gradient, the initial optimality gap and the uniform convexity parameter, and allows inexact solution for each high-order step. To our best knowledge, no parameter-free methods can attain such a fast gradient norm convergence rate which matches that of the function value residual in the convex case, and no such parameter-free methods for uniformly convex problems exist. These results substantially generalize existing parameter-free and inexact high-order methods and recover first-order algorithms as special cases, providing a unified approach for fast gradient minimization across a broad range of smoothness and curvature regimes.
@@ 1 Introduction
The last decade has witnessed an increasing interest in the design and analysis of high-order methods for convex programming. High-order methods exploit high-order derivative information to achieve faster convergence compared to first-order methods. Nesterov and Polyak established the first global worst-case complexity analysis for the second-order schemes by designing a cubic regularization of Newton method (CNM) [30]. They showed that CNM takes at most $\mathcal{O}(\varepsilon^{-1/2})$ iterations to reduce the functional residual below a given precision $\varepsilon,$ i.e., $f(x)-f^{*}\leq\varepsilon,$ where $f$ is a twice-differentiable convex function with a Lipschitz continuous Hessian. Later, Nesterov obtained an improved complexity bound of $\mathcal{O}(\varepsilon^{-1/3})$ by developing an accelerated version of CNM [23]. After that, Monteiro and Svaiter [22] derived an improved iteration complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1})$ through an accelerated Newton proximal extragradient (ANPE). Later, a lower bound of $\Omega(\varepsilon^{-2/7})$ for second-order methods was established [3], which implies ANPE [22] is optimal up to a logarithmic factor. When it comes to $p>2$, the $p$-th order tensor method and its accelerated version are developed for convex functions with Lipschitz continuous $p$-th derivatives (2) in [5, 25]. Specifically, to achieve $f(x)-f^{*}\leq\varepsilon,$ the $p$-th order tensor method takes up to $\mathcal{O}(\varepsilon^{-1/p})$ iterations, while its accelerated version only requires $\mathcal{O}(\varepsilon^{-1/(p+1)})$ iterations [25]. Subsequent methods improved the iteration complexity to $\mathcal{O}(\varepsilon^{-2/(3p+1)}\log\varepsilon^{-1})$, such as those built upon the ANPE method [19, 6, 12], and the inexact high-order proximal-point methods [26, 28]. Recent methods [7, 20] managed to remove the extra logarithmic factor and thus are optimal.
@@
On the other hand, the worst-case function for the class of functions with Lipschitz continuous Hessian has a discontinuous third derivative [28]. Therefore, quite interestingly, some algorithms that rely solely on second-order information can surpass the classical lower bound of $\Omega(\varepsilon^{-2/7})$ for the smaller function class whose second and third derivatives are Lipschitz continuous. For example, a second-order method only requires an iteration complexity of $\mathcal{O}(\varepsilon^{-1/4}),$ at each iteration, it calls the second-order oracle once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most, see Nesterov [27]. Furthermore, in [28], Nesterov shows that a second-order implementation of the third-order accelerated proximal-point method with segment search can achieve an iteration complexity of $\mathcal{O}(\varepsilon^{-1/5}),$ at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most. Following this, second-order methods have achieved the convergence rate of third-order accelerated tensor methods in terms of function value residual for composite problems [2, 1].
@@
Furthermore, the aforementioned methods assume the access to the Lipschitz constant of the corresponding derivative, which is difficult to estimate in practice, and most assume exact solutions for each subproblem step, which is hard to obtain when $p>3$ [25]. Therefore, it is important to develop parameter-free and inexact algorithms that automatically estimate such parameters and solve each step inexactly, while maintaining similar theoretical guarantees. Among the second-order methods, Cartis et al. proposed an adaptive cubic regularized Newton method [8], and further relaxed the criterion for solving each subproblem while maintaining the convergence properties, which is widely used due to its numerical efficiency. Recent works [14, 15, 18, 13, 16] develop accelerated high-order methods that are fully independent of problem constants while maintaining superior theoretical iteration bounds, they achieve an iteration complexity of $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{1}{{p+1}}}}}),$ where $L_{0}$ is the initial line search value and $\theta$ represents the inexactness level of each subproblem.
@@
Despite these progresses of high-order methods for convex optimization, the convergence guarantees in existing works are mainly established for function value residual, i.e., generating a point $\widehat{x}$ such that $f(\widehat{x})-f^{\ast}\leq\varepsilon$, where $f$ is the objective function, $f^{\ast}$ is its optimal value, and $\varepsilon>0$ is a given precision; not on the gradient norm criterion, i.e., $\|\nabla f(\widehat{x})\|\leq\varepsilon.$ While the former is theoretically appealing, it is difficult to verify in practice since $f^{*}$ is unknown. In contrast, the gradient norm criterion is easily computable, may serve as a measure of optimality. Moreover, the gradient norm minimization perspective also provides a practical mechanism to monitor the progress of the algorithm and can be used to estimate problem parameters adaptively. There are some works provided complexity results for the gradient norm criterion; however, they do not match the guarantees for function value residual. For example, in [14], the authors studied a regularized accelerated tensor method and established a complexity of $\mathcal{O}(\varepsilon^{-1/(p+1)}\log\varepsilon^{-1})$, which is worse than the corresponding function value residual complexity of the accelerated tensor methods [25] by a logarithmic factor. As another example, the ANPE method [22], exhibits a gradient norm complexity of $\mathcal{O}(\varepsilon^{-1/3}\log\varepsilon^{-1})$, which is worse than its nearly optimal function value residual complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1}).$ A third example is parameter-free and inexact method, where the best known iteration complexity for gradient norm minimization is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{p+1}{{p(p+2)}}}}})$ [15, 16]. However, this complexity is worse than the function value residual complexity, which is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}\varepsilon^{-\frac{1}{p+1}}).$
@@
To summarize, the convergence rates in terms of gradient norm for different types of high-order methods do not match with the corresponding function value residual. This raises a natural question: is there a unified approach to translate the fast function value residual rate into a matching rate for gradient norm minimization?
@@
To address this question, we propose a unified accumulative regularization (AR) method for gradient minimization and use it to accelerate existing higher-order algorithms, yielding best known gradient norm convergence rates for convex problems. Specifically, for an existing high-order algorithm $\mathcal{A}$, we utilize its fast function value residual decay together with its slow gradient norm decay, and design new gradient norm minimization algorithms that match the convergence rate of the function value residual. It is worth noting that our approach is inspired by the first-order gradient minimization algorithm of Lan et al. [21], and recovers their results when $p=1$. Our contributions can be briefly summarized as follows.
@@
First, for a twice differentiable convex function with a Lipschitz continuous Hessian, we refine the estimating sequence analysis of the accelerated CNM (ACNM) to obtain a slow gradient norm convergence rate of $\mathcal{O}(\varepsilon^{-1/2})$, together with its original fast function residual decrease $\mathcal{O}(\varepsilon^{-1/3})$. Then we design a third-order AR method that uses ACNM as a subroutine. The method restarts intermittently, with each epoch initialized with the output from the previous epoch. Moreover, it applies a single accumulative regularization scheme whose parameter depends on $\varepsilon$ at initialization and then increases geometrically across epochs. We show that, to achieve $\|\nabla f(\widehat{x})\|\leq\varepsilon$, the number of iterations reduces to $\mathcal{O}(\varepsilon^{-1/3})$, thereby matching the complexity for the function value residual.
@@
    Suppose $g$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2,$ by choosing $\mathcal{A}$ as the $p$-th order accelerated tensor method [16], the $(p+\nu)$-th order AR framework only requires $\mathcal{O}(\varepsilon^{-1/{(p+\nu)}})$ iterations, which matches that of the function residual [25].
@@
    Suppose $g$ satisfies (4) with general $p\geq 1,$ and $\nu=1$, by choosing $\mathcal{A}$ as the nearly optimal tensor methods [19, 6, 12] or $p$-th order inexact proximal point method [28, 1], the $(p+1)$-th order AR framework requires $\mathcal{O}\left(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon}\right)$ iterations, which matches that of the function value residual. The extra logarithmic term disappears if $\mathcal{A}$ is chosen as the optimal $p$-th order tensor method (for function-residual) [7, 20] and can output a slow-gradient norm.
@@
    Suppose $g$ satisfies (4) with $p=3,\nu=1,$ by choosing $\mathcal{A}$ as the second-order method [28], the fourth-order AR framework achieves an iteration complexity of $\mathcal{O}(\varepsilon^{-{1}/{5}}),$ at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
Third, we develop inexact and parameter-free high-order AR framework for unconstrained convex optimization (1) where $f$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2.$ By choosing $\mathcal{A}$ as adaptive and inexact tensor methods [13, 16], the $(p+\nu)$-th order inexact and parameter-free AR framework only requires
$\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}\text{dist}(x_{0},X^{*})^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
calls to the $p$-th order oracle, where $L_{0}$ is the initial line search value and $\theta$ represents the inexactness level of the subproblem of each iteration. It does not require the knowledge of the number of iterations to run in advance, the number of restarting epochs, the knowledge of $L_{p,\nu}(f),$ and the initial optimality gap $\text{dist}(x_{0},X^{*})$. It matches the best known convergence rate in terms of function value residual for adaptive and inexact methods [18, 13, 16].
@@
iterations if $q=p+1,$ where $\nu_{0}\in\partial f(x_{0}).$ Notice that when $p=1$, this recovers the gradient minimization results for strongly convex and smooth function [21]. Furthermore, it improves the condition number dependence established in [10, 9] for linear decay of the function value residual. The linear rate improves to super linear rate if $q<p+1,$ i.e., $\|\widehat{\nu}\|\leq\varepsilon$ within at most
@@
iterations. When $p=q=2,$ the algorithm achieves the iterations complexity of $\mathcal{O}({\log\log{\tfrac{\sigma^{2}_{2}(f)}{L_{3}(g)\varepsilon}}}+[{L_{3}(g)\|\nu_{0}\|}/{\sigma^{2}_{2}(f)}]^{{\frac{2}{7}}}),$ which nearly matches the lower bound [3] for function value residual. The algorithm become sublinear if $q>p+1,$ and requires
@@
iterations. It improves over the sublinear rate $\mathcal{O}(\varepsilon^{-\frac{q-p-1}{pq}})$ when $q>p+1$ for function value residual [9].
Lastly, we develop inexact and parameter-free gradient minimization which do not require the knowledge of Lipschitz continuous parameter or uniform convex parameter. Using unconstrained uniformly convex optimization problem with $q=p+1$ as a showcase, our parameter-free and inexact method requires
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right)$
iterations to compute a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$. It is the first parameter-free and inexact method that does not require the knowledge of the uniform convexity parameter $\sigma_{p+1}(f),$ while still matching the fastest known convergence rate of methods that assume $\sigma_{p+1}(f)$ is known. A very recent work [34] derive similar complexity as the prior works [10, 9] and established global and local complexities with linear ($q=p+1$) and superlinear ($q>p+1$) decay by an inexact and adaptive approach. Notice that it still assume the access to the uniform convexity parameter.
To the best of our knowledge, all of the above gradient-norm convergence results are new. The only exception is the recent work [11], which established a complexity of $\mathcal{O}(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon})$ for convex problem. However, the result appears to be more restrictive as they require $p\geq 2,$ thus hard to expand to the cases of bounded Hessian $(p=2,\nu=0)$ or $L$-smooth $(p=1,\nu=1).$ Moreover, it is restricted to the unconstrained case and does not extend to general convex composite or uniformly convex objectives. Unlike AR, it relies on known problem parameters and requires the subproblem of each iteration to be solved exactly.
The rest of this paper is organized as follows. In Section 2, we provide a brief review of the ACNM method and its modified analysis for convex functions with smooth Hessians. Then, we construct a third-order AR approach to accelerate the gradient norm convergence rate. In Section 3, we propose a general AR framework and employ different types of subroutines $\mathcal{A}$ for structured convex composite problems. In Section 4, we derive parameter-free inexact high-order AR framework for convex functions. In Section 5, we further investigate gradient norm minimization for uniformly convex functions and develop a parameter-free and inexact algorithm.

### v2

# Gradient Minimization Using High-Order MethodsYao Ji and Guanghui Lan

@@ Abstract
This paper develops a unified high-order accumulative regularization (AR) framework for convex and uniformly convex gradient norm minimization. Existing high-order methods often exhibit a gap: the function-value residual decreases fast, while the gradient norm converges much slower. To close this gap, we introduce AR that systematically transforms the fast function-value residual convergence rate into a fast (matching) gradient norm convergence rate.
Specifically, for composite convex problems, to compute an approximate solution such that the norm of its (sub)gradient does not exceed $\varepsilon,$ the proposed AR methods match the best corresponding convergence rate for the function-value residual. We further extend the framework to uniformly convex settings, establishing linear, superlinear, and sublinear convergence of the gradient norm under different lower curvature conditions. Moreover, we design parameter-free algorithms that require no input of problem parameters, e.g., the Lipschitz constant of the $p$-th-order gradient, the initial optimality gap and the uniform convexity parameter, and allow an inexact solution for each high-order step. To the best of our knowledge, no parameter-free methods can attain such a fast gradient norm convergence rate which matches that of the function-value residual in the convex case, and no such parameter-free methods for uniformly convex problems exist. These results substantially generalize existing parameter-free and inexact high-order methods and recover first-order algorithms as special cases, providing a unified approach for fast gradient minimization across a broad range of smoothness and curvature regimes.
@@ 1 Introduction
The last decade has witnessed an increasing interest in the design and analysis of high-order methods for convex programming. High-order methods exploit high-order derivative information to achieve faster convergence compared to first-order methods. Nesterov and Polyak established the first global worst-case complexity analysis for the second-order schemes by designing a cubic regularization of Newton method (CNM) [nesterov2006cubic]. They showed that CNM takes at most $\mathcal{O}(\varepsilon^{-1/2})$ iterations to reduce the functional residual below a given precision $\varepsilon,$ i.e., $f(x)-f^{*}\leq\varepsilon,$ where $f$ is a twice-differentiable convex function with a Lipschitz continuous Hessian. Later, Nesterov obtained an improved complexity bound of $\mathcal{O}(\varepsilon^{-1/3})$ by developing an accelerated version of CNM [nesterov2008accelerating]. After that, Monteiro and Svaiter [monteiro2013accelerated] derived an improved iteration complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1})$ through an accelerated Newton proximal extragradient (ANPE). Later, a lower bound of $\Omega(\varepsilon^{-2/7})$ for second-order methods was established [arjevani2019oracle], which implies that ANPE [monteiro2013accelerated] is optimal up to a logarithmic factor. When it comes to $p>2$, the $p$-th-order tensor method and its accelerated version are developed for convex functions with Lipschitz continuous $p$-th derivatives (2) in [baes2009estimate, nesterov2021implementable]. Specifically, to achieve $f(x)-f^{*}\leq\varepsilon,$ the $p$-th-order tensor method takes up to $\mathcal{O}(\varepsilon^{-1/p})$ iterations, while its accelerated version only requires $\mathcal{O}(\varepsilon^{-1/(p+1)})$ iterations [nesterov2021implementable]. Subsequent methods improved the iteration complexity to $\mathcal{O}(\varepsilon^{-2/(3p+1)}\log\varepsilon^{-1})$, such as those built upon the ANPE method [jiang2021optimal, bubeck2019near, gasnikov2019optimal], and the inexact high-order proximal-point methods [nesterov2021inexact, nesterov2023inexact]. Recent methods [carmon2022optimal, kovalev2022first] managed to remove the extra logarithmic factor and thus are optimal.
@@
On the other hand, the worst-case function for the class of functions with Lipschitz continuous Hessian has a discontinuous third derivative [nesterov2023inexact]. Therefore, quite interestingly, some algorithms that rely solely on second-order information can surpass the classical lower bound of $\Omega(\varepsilon^{-2/7})$ for the smaller function class whose second and third derivatives are Lipschitz continuous. For example, a second-order method only requires an iteration complexity of $\mathcal{O}(\varepsilon^{-1/4});$ at each iteration, it calls the second-order oracle once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most, see Nesterov [nesterov2021superfast]. Furthermore, in [nesterov2023inexact], Nesterov shows that a second-order implementation of the third-order accelerated proximal-point method with segment search can achieve an iteration complexity of $\mathcal{O}(\varepsilon^{-1/5}),$ at each iteration, it calls the second-order oracle once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most. Following this, second-order methods have achieved the convergence rate of third-order accelerated tensor methods in terms of function-value residual for composite problems [ahookhosh2024high, ahookhosh2021high].
@@
Furthermore, the aforementioned methods assume access to the Lipschitz constant of the corresponding derivative, which is difficult to estimate in practice, and most assume exact solutions for each subproblem step, which is hard to obtain when $p>3$ [nesterov2021implementable]. Therefore, it is important to develop parameter-free and inexact algorithms that automatically estimate such parameters and solve each step inexactly, while maintaining similar theoretical guarantees. Among the second-order methods, Cartis et al. proposed an adaptive cubic regularized Newton method [cartis2012evaluation], and further relaxed the criterion for solving each subproblem while maintaining the convergence properties, which is widely used due to its numerical efficiency. Recent works [grapiglia2020tensor, grapiglia2022tensor, jiang2020unified, grapiglia2023adaptive, He2025HAR] develop accelerated high-order methods that are fully independent of problem constants while maintaining superior theoretical iteration bounds; they achieve an iteration complexity of $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{1}{{p+1}}}}}),$ where $L_{0}$ is the initial line-search value and $\theta$ represents the inexactness level of each subproblem.
@@
Despite these advances of high-order methods for convex optimization, the convergence guarantees in existing works are mainly established for the function-value residual, i.e., generating a point $\widehat{x}$ such that $f(\widehat{x})-f^{\ast}\leq\varepsilon$, where $f$ is the objective function, $f^{\ast}$ is its optimal value, and $\varepsilon>0$ is a given precision; but not on the gradient norm criterion, i.e., $\|\nabla f(\widehat{x})\|\leq\varepsilon.$ While the former is theoretically appealing, it is difficult to verify in practice since $f^{*}$ is unknown. In contrast, the gradient norm criterion is easily computable and may serve as a measure of optimality. Moreover, the gradient norm minimization perspective also provides a practical mechanism to monitor the progress of the algorithm and can be used to estimate problem parameters adaptively. Some works provide complexity results for the gradient norm criterion; however, they do not match the guarantees for the function-value residual. For example, in [grapiglia2020tensor], the authors studied a regularized accelerated tensor method and established a complexity of $\mathcal{O}(\varepsilon^{-1/(p+1)}\log\varepsilon^{-1})$, which is worse than the corresponding function-value residual complexity of the accelerated tensor methods [nesterov2021implementable] by a logarithmic factor. As another example, the ANPE method [monteiro2013accelerated] exhibits a gradient norm complexity of $\mathcal{O}(\varepsilon^{-1/3}\log\varepsilon^{-1})$, which is worse than its nearly optimal function-value residual complexity of $\mathcal{O}(\varepsilon^{-2/7}\log\varepsilon^{-1}).$ A third example is a parameter-free and inexact method, where the best-known iteration complexity for gradient norm minimization is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}{\varepsilon^{-{\frac{p+1}{{p(p+2)}}}}})$ [grapiglia2022tensor, grapiglia2023adaptive]. However, this complexity is worse than the function-value residual complexity, which is $\mathcal{O}(\max\{L_{0},pL_{p+1}(f),\theta\}\varepsilon^{-\frac{1}{p+1}}).$
@@
To summarize, the convergence rates in terms of gradient norm for different types of high-order methods do not match the corresponding function-value residual. This raises a natural question: is there a unified approach to translate the fast function-value residual rate into a matching rate for gradient norm minimization?
@@
To address this question, we propose a unified accumulative regularization (AR) method for gradient minimization and use it to accelerate existing high-order algorithms, yielding best-known gradient norm convergence rates for convex problems. Specifically, for an existing high-order algorithm $\mathcal{A}$, we utilize its fast function-value residual decay together with its slow gradient norm decay, and design new gradient norm minimization algorithms that match the convergence rate of the function-value residual. It is worth noting that our approach is inspired by the first-order gradient minimization algorithm of Lan et al. [lan2024optimalparameterfreegradientminimization], and recovers their results when $p=1$. Our contributions can be briefly summarized as follows.
@@
First, for a twice-differentiable convex function with a Lipschitz continuous Hessian, we refine the estimating sequence analysis of the accelerated CNM (ACNM) to obtain a slow gradient norm convergence rate of $\mathcal{O}(\varepsilon^{-1/2})$, together with its original fast function-value residual decrease $\mathcal{O}(\varepsilon^{-1/3})$. Then we design a third-order AR method that uses ACNM as a subroutine. The method restarts intermittently, with each epoch initialized with the output from the previous epoch. Moreover, it applies a single accumulative regularization scheme whose parameter depends on $\varepsilon$ at initialization and then increases geometrically across epochs. We show that, to achieve $\|\nabla f(\widehat{x})\|\leq\varepsilon$, the number of iterations reduces to $\mathcal{O}(\varepsilon^{-1/3})$, thereby matching the complexity for the function-value residual.
@@
    Suppose $g$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2,$ by choosing $\mathcal{A}$ as the $p$-th-order accelerated tensor method [grapiglia2023adaptive], the $(p+\nu)$-th order AR framework only requires $\mathcal{O}(\varepsilon^{-1/{(p+\nu)}})$ iterations, which matches that of the function-value residual [nesterov2021implementable].
@@
    Suppose $g$ satisfies (4) with general $p\geq 1,$ and $\nu=1$, by choosing $\mathcal{A}$ as the nearly optimal tensor methods [jiang2021optimal, bubeck2019near, gasnikov2019optimal] or $p$-th-order inexact proximal point method [nesterov2023inexact, ahookhosh2021high], the $(p+1)$-th order AR framework requires $\mathcal{O}\left(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon}\right)$ iterations, which matches that of the function-value residual. The extra logarithmic term disappears if $\mathcal{A}$ is chosen as the optimal $p$-th-order tensor method (for function-residual) [carmon2022optimal, kovalev2022first] and can output a slow-gradient norm.
@@
    Suppose $g$ satisfies (4) with $p=3,\nu=1,$ by choosing $\mathcal{A}$ as the second-order method [nesterov2023inexact], the fourth-order AR framework achieves an iteration complexity of $\mathcal{O}(\varepsilon^{-{1}/{5}});$ at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
Third, we develop an inexact and parameter-free high-order AR framework for unconstrained convex optimization (1) where $f$ satisfies (4) with general $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2.$ By choosing $\mathcal{A}$ as adaptive and inexact tensor methods [grapiglia2020tensor, grapiglia2023adaptive], the $(p+\nu)$-th order inexact and parameter-free AR framework only requires
$\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}\operatorname{dist}(x_{0},X^{*})^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
calls to the $p$-th-order oracle, where $L_{0}$ is the initial line-search value and $\theta$ represents the inexactness level of the subproblem of each iteration. It does not require knowledge of the number of iterations to run in advance, the number of restarting epochs, or $L_{p,\nu}(f),$ and the initial optimality gap $\operatorname{dist}(x_{0},X^{*})$. It matches the best-known convergence rate in terms of function-value residual for adaptive and inexact methods [jiang2020unified, grapiglia2020tensor, grapiglia2023adaptive, He2025HAR].
@@
iterations if $q=p+1,$ where $\nu_{0}\in\partial f(x_{0}).$ Notice that when $p=1$, this recovers the gradient minimization results for strongly convex and smooth function [lan2024optimalparameterfreegradientminimization]. Furthermore, it improves the condition number dependence established in [doikov2022local, cartis2022evaluation] for linear decay of the function-value residual. The linear rate improves to superlinear rate if $q<p+1,$ i.e., $\|\widehat{\nu}\|\leq\varepsilon$ within at most
@@
iterations. When $p=q=2,$ the algorithm achieves the iteration complexity of $\mathcal{O}({\log\log{\tfrac{\sigma^{2}_{2}(f)}{L_{3}(g)\varepsilon}}}+[{L_{3}(g)\|\nu_{0}\|}/{\sigma^{2}_{2}(f)}]^{{\frac{2}{7}}}),$ which nearly matches the lower bound [arjevani2019oracle] for the function-value residual. The algorithm becomes sublinear if $q>p+1,$ and requires
@@
iterations. It improves over the sublinear rate $\mathcal{O}(\varepsilon^{-\frac{q-p-1}{pq}})$ when $q>p+1$ for the function-value residual [cartis2022evaluation]. The results in [doikov2022local] are non-asymptotic rates, but assume knowledge of the Lipschitz constant $L_{p+1}(q)$ and uniform convexity $\sigma_{q}(f)$; a very recent work [Welzel2025Local] removes the need to know the Lipschitz constant $L_{p+1}(q)$ through an adaptive algorithm, but it is an asymptotic rate.
Lastly, we develop inexact and parameter-free gradient minimization which does not require knowledge of the Lipschitz continuous parameter or the uniform convex parameter. Using the unconstrained uniformly convex optimization problem with $q=p+1$ as a showcase, our parameter-free and inexact method requires
$\mathcal{O}\left(\left(\tfrac{\max\{{pL_{p+1}}(f),L_{0},\theta\}}{\sigma_{p+1}(f)}\right)^{\frac{1}{p+1}}\left\{\left\lceil\log_{4}\tfrac{\sigma_{p+1,0}}{\sigma_{p+1}(f)}\right\rceil+\left\lceil\log_{2}\tfrac{\|\nabla f(x_{0})\|}{\varepsilon}\right\rceil\right\}\right)$
iterations to compute a solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon,$ where $\sigma_{p+1,0}$ is the initial guess of the uniform convexity parameter $\sigma_{p+1}(f).$ It is the first parameter-free and inexact method that does not require the knowledge of the uniform convexity parameter $\sigma_{p+1}(f),$ while still matching the fastest known non-asymptotic convergence rate of methods that assume $L_{p+1}(f),\sigma_{p+1}(f)$ are known.
To the best of our knowledge, all of the above gradient norm convergence results are new. The only exception is the recent work [dvurechensky2024near], which established a complexity of $\mathcal{O}(\varepsilon^{-\frac{2}{3p+1}}\log\tfrac{1}{\varepsilon})$ for convex problems. However, the result appears to be more restrictive as it requires $p\geq 2,$ thus hard to expand to the cases of bounded Hessian $(p=2,\nu=0)$ or $L$-smooth $(p=1,\nu=1).$ Moreover, it is restricted to the unconstrained case and does not extend to general convex composite or uniformly convex objectives. Unlike AR, it relies on known problem parameters and requires the subproblem in each iteration to be solved exactly.
The rest of this paper is organized as follows. In Section 2, we provide a brief review of the ACNM method and its modified analysis for convex functions with smooth Hessians. Then, we construct a third-order AR approach to accelerate the gradient norm convergence rate. In Section 3, we propose a general AR framework and employ different types of subroutines $\mathcal{A}$ for structured convex composite problems. In Section 4, we derive a parameter-free, inexact, high-order AR framework for convex functions. In Section 5, we further investigate gradient norm minimization for uniformly convex functions and develop a parameter-free and inexact algorithm.